What is x-4y+20=0 in slope-intercept form?
Answer by Ultimatemagic(23)

Y=x+5
The first step is to get x on the other side so x is subtracted from both sides, so far the problem should look like this -4y+20=x-0, now to get y by its self you minus 20 from both sides, now it should look like this (-4y=x-20) now to officially get y by its self divide -4 by y and -20 which will dive you y=x+5